# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v2

@@ Abstract
In deep Reinforcement Learning (RL) models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence. Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps. To adapt the learning rate, a standard technique consists of using decay schedulers. However, these schedulers assume that the model is progressively approaching convergence, which may not always be true, leading to delayed or premature adjustments. In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that selects the learning rate based on the agent’s performance during training. LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution. Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms for some tasks.
@@ 1 Introduction
Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.37, 20, 43, 44). While ANNs were previously used as value function approximators [32], the introduction of Deep Q-Networks (DQN) by Mnih et al. [25, 26] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
Experience replay (see 23) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation. The target network further stabilizes the learning process by periodically copying the parameters of the learning network. This strategy is necessary because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence. By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process. Additionally, the learning rate [^1] The terms “learning rate” and “step-size” are used interchangeably in the literature and they technically refer to the same concept. hyperparameter controls the magnitude of these gradient steps in optimizers such as the Stochastic Gradient Descent (SGD) algorithm, affecting the training convergence.
The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3]. Determining the appropriate learning rate is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 15, 7, 46). However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing. In order to adapt its initial choice during training, optimizers such as Adam [18] and RMSProp [41] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information. Therefore, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [35, 46]. Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge). As the policy improves, the data distribution encountered by the agent becomes more concentrated, but this evolution occurs at a different pace than the overall training progress. For instance, some environments require extensive exploration due to the sparseness of rewards, while others need more exploitation to refine the policy to the complexity of the task. Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
Contributions.
In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL. Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance. Our method has the advantage of being algorithm-agnostic and applicable to any optimizer. By exploiting different configurations, we aim to illustrate the use cases of LRRL and assess its robustness. We provide experiments with the Adam and RMSProp optimizers, with different sets of arms, and two deep RL algorithms (DQN and IQN). Tests are carried out across a variety of Atari games, using baselines provided in the Dopamine framework (see 9). We also show the effectiveness of LRRL using a standard SGD optimizer in a non-RL optimization context. Related work is detailed in Appendix A. Our main contributions and results can be summarized as follows:
  • We introduce LRRL, a novel approach that leverages a MAB algorithm to dynamically select the learning rate during training in deep RL.
  • We assess LRRL robustness across a set of Atari games, using different configurations, optimizers and deep RL algorithms. Our empirical results demonstrate that LRRL achieves performance competitive with or superior to standard learning using fixed learning rates or decay schedulers. We further highlight its advantages by considering an SGD optimizer on stationary non-convex landscapes.
  • More generally, we show that LRRL matches the scheduler’s behavior by dynamically choosing lower learning rate values as the training progresses.
@@ 2 Preliminaries
This section introduces the RL and MAB frameworks, defining supporting notation.
@@ 2.1 Deep Reinforcement Learning
An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.39 for details). In RL, starting from an initial state $s_{0}$, a learner called the agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$. In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(\cdot|s_{t},a_{t})$. The agent’s objective is to learn a policy $\pi(\cdot|s)$ that maps a state $s$ to a probability distribution over actions in $\mathcal{A}$, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
To learn how to perform a task, value function-based algorithms coupled with ANNs [25, 26] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(\cdot|s)=1_{a^{*}(s)}(\cdot)$ where $1_{a^{*}(s)}$ denotes the uniform distribution over ${a^{*}(s)}$ with $a^{*}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$. By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
@@
where $\theta^{-}$ are the target network parameters used to compute the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$. The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$. The minimization, and hence the update of the parameters $\theta$, is done according to the optimizer’s routine. A simple possibility is to use SGD using mini-batch approximations of the loss gradient:
@@
with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate. Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time. This continuous evolution introduces instability during learning, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$. However, stability also depends on how the parameters $\theta$ change during each update. This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ over the training steps.
@@ 2.2 Multi-Armed Bandits
MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 21). MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$. Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
@@
where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the (pontentially changing) single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm. A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
@@ 3 Dynamic Learning Rate for Deep RL
In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic LRRL. LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance. LRRL couples with SGD optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm. As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
@@ 3.1 Selecting the Learning Rate Dynamically
Our problem can be framed as selecting a learning rate $\eta$ for policy updates —specifically, when updating the parameters $\theta$ after $\lambda$ interactions with the environment.
Before training, a set $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ of $K$ learning rates are defined by the user. Then, during training, a MAB algorithm selects, at every round $n$ —that is, at every $\kappa$ interactions with the environment— an arm $k_{n}\in\{1,\dots,K\}$ according to a probability distribution $p_{n}$ defined based on previous rewards, as explained in the next section. The parameters $\theta$ are then updated using the sampled learning rate $\eta_{k_{n}}$. The steps involved in this meta-learning approach are summarized in Algorithm 1.
@@
Parameters:
    Set of learning rates $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$
    Number of episodes $M$
    Horizon length $T$
    Update window $\lambda$ for the learning network $\theta$
    Update window $\tau$ for the target network $\theta^{-}$
    Update window $\kappa$ for arm probabilities $p$
Initialize:
    Parameters $\theta$ and $\theta^{-}$
    Arm probabilities $p_{0}\leftarrow(\tfrac{1}{K},\dots,\tfrac{1}{K})$
    MAB round $n\leftarrow 0$
    Cumulative reward $R\leftarrow 0$
    Environment interactions counter $C\leftarrow 0$
@@
         Choose action $a_{t}$ following the policy $\pi_{\theta}(s)$ with probability $1-\epsilon$triangleright $(\epsilon$-greedy strategy)
@@
         Increase environment interactions counter $C\leftarrow C+1$
@@
                  Compute average of the last $C$ rewards $f_{n}\leftarrow\frac{R}{C}$
@@
                  Reset $R\leftarrow 0$ and $C\leftarrow 0$
@@
Note that the same algorithm might be used with learning rates schedulers, that is with $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ where $\eta_{k}:\mathbb{N}\to\mathbb{R}_{+}$ is a predefined function, usually converging towards 00 at infinity. If so, the learning rate used at round $n$ of the optimization is $\eta_{k_{n}}(n)$.
@@ 3.2 Updating the Probability Distribution
As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback. To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction). At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively. Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to changes of the best action and to improvements in policy performance.
Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions. Based on Moskovitz et al. [27], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
@@
The improvement in performance allows computation of the next weights $w_{n+1}$ as follows, where initially $w_{1}=(0,\ldots,0)$:
$\forall k\in\{1,\ldots,K\},\quad w_{n+1}(k)=\begin{cases}\delta\,w_{n}(k)+\alpha\,\frac{f^{\prime}_{n}}{e^{w_{n}(k)}}&\text{if }k=k_{n}\\ \delta\,w_{n}(k)&\text{otherwise}\,,\end{cases}$ (3)
where $\alpha>0$ is a step-size parameter. The distribution $p_{n+1}$, used to draw the next arm $k_{n+1}$, is

### v3

@@ Abstract
In deep Reinforcement Learning (RL), the learning rate critically influences both stability and performance, yet its optimal value shifts during training as the environment and policy evolve. Standard decay schedulers assume monotonic convergence and often misalign with these dynamics, leading to premature or delayed adjustments. We introduce LRRL, a meta-learning approach that dynamically selects the learning rate based on policy performance rather than training steps. LRRL adaptively favors rates that improve returns, remaining robust even when the candidate set includes values that individually cause divergence. Across Atari and MuJoCo benchmarks, LRRL achieves performance competitive with or superior to tuned baselines and standard schedulers. Our findings position LRRL as a practical solution for adapting to non-stationary objectives in deep RL.
@@ 1 Introduction
Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.39, 21, 47, 48). These achievements largely depend on training neural networks with gradient-based optimizers, where the choice of learning rate is critical to stability and convergence. An overly large learning rate can lead to divergence, while a small one can result in prohibitively slow learning (see 17, 6, 50). Despite its importance, selecting an appropriate learning rate in RL remains challenging due to two intertwined factors: the non-stationarity of the objective and the high cost of hyperparameter tuning.
Unlike supervised learning, where the loss function is typically stationary, RL agents learn from data generated by an evolving policy. As training progresses, both the distribution of visited states and the policy-induced rewards change, making the optimization landscape highly dynamic. Consequently, a learning rate that works well early in training may later become suboptimal. Prior work shows that even simple adjustments, such as reducing the learning rate in later stages, can improve performance in deep RL [2]. However, designing an effective schedule is difficult because training steps do not necessarily reflect monotonic convergence toward a better policy.
Decay schedules, such as linear or exponential decay [37, 50], reduce the learning rate over time based on predefined rules, assuming that the optimization problem becomes easier as training proceeds. This assumption often fails in RL, where progress can plateau or regress due to exploration challenges, sparse rewards, or environment stochasticity. Similarly, adaptive optimizers such as RMSProp [45] and Adam [20] adjust effective learning rates (see 25) based on past gradient statistics, which stabilizes updates but does not account for shifts in policy quality. Because these methods rely on step counts or gradient norms rather than environment feedback, they often make premature or delayed adjustments, limiting their ability to cope with RL’s non-stationarity.
This motivates the question: Can we adapt the learning rate dynamically based on the agent’s performance, rather than relying on training progress or gradient-based heuristics? A method satisfying this criterion should (i) leverage feedback that reflects policy quality rather than indirect signals like gradient norms, (ii) adapt to the different dynamics of early exploration and late-stage refinement, and (iii) remain general enough to integrate with a wide range of algorithms and optimizers.
We address this challenge by proposing Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that casts learning rate adaptation as a dual objective: while the RL agent seeks to maximize cumulative returns, the bandit algorithm minimizes regret in selecting the most effective learning rates over time. LRRL formulates this as an adversarial multi-armed bandit (MAB) problem, where each arm corresponds to a candidate learning rate, and feedback is based on policy performance. By leveraging adversarial bandit algorithms with time-decay weighting, LRRL balances exploration of alternatives with exploitation of promising rates. Unlike fixed schedules or gradient-based meta-learning heuristics, LRRL is algorithm-agnostic, seamlessly integrates with optimizers such as Adam, RMSProp, and SGD, and enables dynamic adaptation across different phases of training.
We evaluate LRRL across a diverse set of scenarios, including discrete and continuous control tasks (Atari and MuJoCo), multiple deep RL algorithms (DQN, IQN, and PPO), and various optimizers. Our results show that LRRL achieves competitive or superior performance compared to fixed learning rates and standard decay strategies, while also demonstrating robustness to variations in configuration. Furthermore, to isolate RL-specific confounders, such as exploration and reward sparsity, we assess LRRL on stationary non-convex optimization problems using SGD, illustrating its ability to adaptively combine aggressive and conservative updates in a single optimization trajectory. Our main contributions are:
  • We introduce LRRL, a novel approach that leverages a bandit-based controller to select among a small set of learning rates on the fly using policy performance as feedback. This directly targets RL’s non-stationarity and avoids assuming monotonic convergence.
  • We demonstrate robustness to poorly performing candidates: LRRL remains stable and competitive even when the candidate set contains rates that individually fail, by adapting toward effective rates during training.
  • We evaluate LRRL using a set of learning algorithms on Atari and MuJoCo tasks, showing performance competitive with or superior to tuned fixed-rate training and decay schedulers, alongside ablations of key design parameters.
  • Beyond RL, we illustrate how combining multiple rates within a single run can be beneficial in stationary non-convex optimization with SGD, underscoring LRRL’s generality.
@@ 2 Preliminaries
This section introduces the RL and MAB frameworks and defines the supporting notation.
@@ 2.1 Deep Reinforcement Learning
An RL task is defined by a Markov Decision Process (MDP), that is, by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see e.g.42 for details). In RL, starting from an initial state $s_{0}$, a learner called agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$. In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(\cdot|s_{t},a_{t})$. The agent’s objective is to learn a policy $\pi(\cdot|s)$ that maps a state $s$ to a probability distribution over actions in $\mathcal{A}$, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl[\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr]$.
To learn how to perform a task, value function-based algorithms coupled with ANNs [28, 29] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(\cdot|s)=1_{a^{*}(s)}(\cdot)$ where $1_{a^{*}(s)}$ denotes the uniform distribution over $a^{*}(s)$} with $a^{*}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$. By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ in replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
@@
where $\theta^{-}$ are the parameters of the target network used to calculate the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$. The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$. The minimization, and hence the update of the parameters $\theta$, is done according to the optimizer’s routine. A simple possibility is to use SGD using mini-batch approximations of the loss gradient:
@@
with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called learning rate. Unlike supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time. This continuous evolution introduces instability during learning, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$. However, stability also depends on how the parameters $\theta$ change during each update. This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ in the training steps.
@@ 2.2 Multi-Armed Bandits
MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 22). MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$. Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
@@
where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the potentially changing single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm. A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$. This guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
@@ 3 Dynamic Learning Rate for Deep RL
In this section, we introduce LRRL, a meta-learning approach for dynamically adapting the learning rate based on policy performance. LRRL addresses the challenge of non-stationarity in RL by framing learning rate selection as a multi-armed bandit (MAB) problem. While the RL policy aims to maximize cumulative return, the bandit algorithm minimizes regret in identifying the most effective learning rates over time. Each candidate learning rate corresponds to an arm, and the bandit uses policy performance as feedback. We adopt an adversarial bandit algorithm with time-decay weighting to accommodate the evolving reward distribution as the agent interacts with the environment.
@@ 3.1 Selecting the Learning Rate Dynamically
Our problem can be framed as selecting a learning rate $\eta$ for policy updates —specifically, when updating the parameters $\theta$ after $\lambda$ interactions with the environment. Before training, the user defines a set $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ of $K$ learning rates. Then, during training, a MAB algorithm selects, at every round $n$ —that is, at every $\kappa$ interactions with the environment— an arm $k_{n}\in\{1,\dots,K\}$ according to a probability distribution $p_{n}$ defined based on previous rewards, as explained in the next section. The parameters $\theta$ are then updated using the sampled learning rate $\eta_{k_{n}}$. The steps involved in this meta-learning approach are summarized in Algorithm 1.
@@
Initialize: network parameters $\theta$, target parameters $\theta^{-}$, arm probabilities $p_{0}\leftarrow(\tfrac{1}{K},\dots,\tfrac{1}{K})$, MAB round $n\leftarrow 0$, cumulative reward $R\leftarrow 0$, counter $C\leftarrow 0$
@@
   Choose action $a_{t}$ following $\pi_{\theta}(\cdot|s)$ with probability $1-\epsilon$   $(\epsilon$-greedy strategy)
@@
   Increase interactions counter $C\leftarrow C+1$
@@
      Compute average $f_{n}\leftarrow\frac{R}{C}$
@@
      Reset $R\leftarrow 0$, $C\leftarrow 0$
@@
Note that the same algorithm might be used with learning rates schedulers, that is with $\mathcal{K}=\{\eta_{1},\dots,\eta_{K}\}$ where $\eta_{k}:\mathbb{N}\to\mathbb{R}_{+}$ is a predefined function, usually converging towards 0 at infinity. If so, the learning rate used at round $n$ of the optimization is $\eta_{k_{n}}(n)$.
@@ 3.2 Updating the Probability Distribution
Since the agent’s performance—and thus the cumulative reward—changes over time, the MAB algorithm naturally receives non-stationary feedback. To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 5 for an introduction). At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively. Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to changes of the best action and to improvements in policy performance. Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions. Based on Moskovitz et al. [30], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
@@
The improvement in performance allows the computation of the next weights $w_{n+1}$ as follows:
[…]
where initially $w_{1}=(0,\ldots,0)$ and $\alpha>0$ is a step-size parameter. The distribution $p_{n+1}$, used to draw the next arm $k_{n+1}$, is defined by: